OpenStudy (anonymous):

Imagine you each need to apply for an unsubsidized loan to pay for the costs of college. The loan has a fixed interest rate of 6.8% compounded monthly. The interest begins immediately, but payments are not required until six months after graduation. The loan must be paid off within 10 years after the deferment period. You must show your work on all of the following calculations, except where indicated. 1. Assuming it takes four years to graduate and no payments are made during the deferment period, what is the balance of the loan when it’s time to make the first payment?

OpenStudy (anonymous):

For number 1 my answer was B= 17,426(1+0.68/12)^10 B=17,426(1.005666667)^10 B=30,239.72

OpenStudy (anonymous):

So what I need help with is figuring out this question What monthly payment must be made to pay off the loan within 10 years?

OpenStudy (amistre64):

\[B_{10}=B(k)^{10}-P\frac{1-k^{10}}{1-k}\] assuming the balance at the end is zero: \[0=B(k)^{10}-P\frac{1-k^{10}}{1-k}\]solve for P

OpenStudy (amistre64):

might need to adjust the exponents to match the number of periods tho .... 10*12

OpenStudy (anonymous):

O___________O

OpenStudy (anonymous):

I've never seen that equation in my life

OpenStudy (amistre64):

i use it all the time ....

OpenStudy (amistre64):

and your #1 is off

OpenStudy (amistre64):

we can use the setup with the original balance if we adjust the exponent by the defered period by adding on 4.5*12 periods

OpenStudy (anonymous):